# Calibrating Large Language Models with Sample Consistency

### v1

@@ Abstract
Accurately gauging the confidence level of Large Language Models’ (LLMs) predictions is pivotal for their reliable application. However, LLMs are often uncalibrated inherently and elude conventional calibration techniques due to their proprietary nature and massive scale. In this work, we explore the potential of deriving confidence from the distribution of multiple randomly sampled model generations, via three measures of consistency. We perform an extensive evaluation across various open and closed-source models on nine reasoning datasets. Results show that consistency-based calibration methods outperform existing post-hoc approaches. Meanwhile, we find that factors such as intermediate explanations, model scaling, and larger sample sizes enhance calibration, while instruction-tuning makes calibration more difficult. Moreover, confidence scores obtained from consistency have the potential to enhance model performance. Finally, we offer practical guidance on choosing suitable consistency metrics for calibration, tailored to the characteristics of various LMs. *** Equal contribution; Qing Lyu did her work while interning at Allen Institute for AI.
@@
Qing Lyu ${}^{*\spadesuit}$  Kumar Shridhar ${}^{*\blacklozenge}$   Chaitanya Malaviya ${}^{\spadesuit}$  Li Zhang ${}^{\spadesuit}$  Yanai Elazar ${}^{\clubsuit}$$\varheartsuit\varheartsuit{}^{\varheartsuit}$ Niket Tandon ${}^{\clubsuit}$  Marianna Apidianaki ${}^{\spadesuit}$  Mrinmaya Sachan ${}^{\blacklozenge}$  Chris Callison-Burch ${}^{\spadesuit}$${}^{\clubsuit}$$^{\spadesuit}$ University of Pennsylvania  ${}^{\blacklozenge}$ ETH Zurich $\varheartsuit\varheartsuit{}^{\varheartsuit}$ University of Washington  ${}^{\clubsuit}$ Allen Institute for AI lyuqing@sas.upenn.edushkumar@ethz.ch
@@ 3 Method
Consistency over multiple generations can be used as an indicator for understanding the confidence associated with model predictions. It has been studied in the past for logit-based uncertainty estimation such as model ensembling Lakshminarayanan et al. (2017) and we extend it to multiple generations in LLMs. For a given input $x$, we generate a set of $n$ candidate outputs $\hat{s}_{1},\dots\hat{s}_{n}$. From each sample $\hat{s}_{i}$, we parse the final answer $\hat{a_{i}}$ using regex matching. We do a majority voting over the entire answer (multi-)set $\mathbf{a}=\{\hat{a_{1}}\dots\hat{a_{n}}\}$ to get the most-voted answer $\bar{a}=\operatorname*{arg\,max}_{a}\sum\nolimits_{i=1}^{n}\mathbbm{1}(\hat{a_{i}}=a)$, where $a$ takes on values from the set of unique answers $\mathbf{\bar{a}}$.
@@ 3.1 Calibration with Consistency
$\texttt{Ent}(\mathbf{a})=1-(-\frac{1}{\log(|\mathbf{\bar{a}}|)}\sum\limits_{i=1}^{|\mathbf{\bar{a}}|}\ p_{i}\ \log(p_{i}))$ (2)
@@ 4 Experimental Setup
$ECE=\sum_{m=1}^{M}\frac{\lvert B_{m}\rvert}{N}\lvert\texttt{acc}(B_{m})-\texttt{conf}(B_{m})\rvert$ (5)
@@
$\texttt{acc}(B_{m})$ $=\frac{1}{\lvert B_{m}\rvert}\sum_{x_{j}\in B_{m}}\mathbb{I}(\hat{y}_{j}=y_{j})$ (6)
$\texttt{conf}(B_{m})$ $=\frac{1}{\lvert B_{m}\rvert}\sum_{x_{j}\in B_{m}}\texttt{conf}(x_{j},\hat{y}_{j})$ (7)
@@
Codex
.175
.151
.159
.249
.249
.209
.188
.179
GPT-3.5-turbo
.205
.221†
.207
.271
.273
n/a
n/a
n/a
GPT-4
.116
.119†
.114
.154
.181
n/a
n/a
n/a
Table 1: Consistency metrics result in better Brier Scores than baselines ($\downarrow$) for closed-source models. Scores are averaged across four domains and five prompting strategies. The best scores are in bold and the second-best scores are underlined. † indicates that the consistency metric performs statistically significantly better than the best-performing baseline ($p<0.05$).
@@
LLaMA-7B
.241†
.232
.235
.474
LLaMA-13B
.222†
.204
.211
.389
LLaMA-70B
.182†
.154
.165
.252
Mistral-7B
.205†
.183
.191
.324
Mistral-7B-instruct
.220†
.216
.215
.384
Table 2: Consistency metrics result in better Brier Scores ($\downarrow$) than the logit baseline for open-source models.
@@
We consider the following LLMs: LLaMA (7B/13B/70B), Mistral (7B/7B-instruct), Codex, GPT-3.5-turbo, and GPT-4.444See checkpoint names and computational resources in Appendix B.
@@ 5.1 Comparing Calibration Methods
Figure 3: Brier Scores ($\downarrow$) are improved with explanation-based prompting strategies, with Chain of Thought (CoT) and Faithful CoT (FCoT) performing the best. Scores here are averaged across all datasets and consistency metrics.
@@
Figure 4: The Brier Score ($\downarrow$) tends to improve as the model size increases for the 3 studied calibration metrics across most of the prompting techniques we consider.
@@ 6.2 How Does Instruction-Tuning Affect Calibration?
Figure 5: Surprisingly, the non-instruction-tuned model (Mistral-7B) has better Brier Scores ($\downarrow$) compared to an instruction-tuned model (Mistral-7B-instruct) across nearly all of our prompting strategies and tasks.
Figure 6: Brier Scores ($\downarrow$) improve as we increase the number of samples for 3 of the 4 datasets. Results are obtained with GPT-3.5-turbo and CoT prompting.
Figure 7: Left: Consistency-based calibration outperforms the verbalized baselines in their ability to discriminate the correctness of predictions measured by Macro-F1 ($\uparrow$). Right: Consistency-based calibration leads to a larger improvement in answer accuracy ($\uparrow$) after correcting the top-kmost uncertain predictions with oracle answers. Scores in these figures are averaged across all prompting strategies.
@@ 6.3 How does the Number of Generated Outputs impact Calibration?
We analyze the usefulness of consistency-based calibration by generating different numbers of output samples and calculating different consistency metrics over them. Figure 6 demonstrates that generating more samples can lead to better calibration scores, as indicated by a downward trend in the Brier Score. We observe the improvement in the Brier Score as a function of the number of samples and the decision of the appropriate number of samples can be made based on the available computational budget and the desired calibration properties. Brier Scores usually saturate after $15-20$ samples, with a sharp drop at the beginning. For budget constraints, $3-5$ samples is a good choice.
@@ 7 Case Study: Does Calibration Help Improve Model Performance?
The second experiment assesses the impact of a calibration method on answer accuracy, assuming an oracle subsequent self-correction mechanism. Given a model’s predictions $\hat{Y}$ on a dataset $X$, we identify the top-kmost uncertain predictions, $\hat{Y_{-}}$, which are those with the lowest confidence scores according to the calibration method, as incorrect. This fixed $k$ is chosen to be the true error rate of all model predictions, i.e., $k=1-\text{acc}(\hat{Y},X)$. Finally, we correct $\hat{Y_{-}}$ with the ground-truth answers and evaluate the resulting accuracy. As shown in Figure 7 (right), post-correction accuracy exceeds original accuracy to the greatest extent when applying consistency-based calibration.
@@ Appendix D Dataset Details
Math Word Problems
GSM8K
8
1,319
Q: Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May?
@@
SVAMP
8
1,000
Q: Each pack of dvds costs 76 dollars. If there is a discount of 25 dollars on each pack. How much do you have to pay to buy each pack?
@@
MultiArith
8
600
Q: For Halloween Debby and her sister combined the candy they received. Debby had 32 pieces of candy while her sister had 42. If they ate 35 pieces the first night, how many pieces do they have left?
@@
ASDiv
8
2,096
Q: Seven red apples and two green apples are in the basket. How many apples are in the basket?
@@
StrategyQA
6
2,290
Q: Did Aristotle use a laptop?
@@
10
359
Q: Yesterday was April 30, 2021. What is the date tomorrow in MM/DD/YYYY?
@@
10
977
Q: Is the following sentence plausible: “Lebron James hit the turnaround jumper”?
@@
Planning
SayCan
7
103
Q: Could you get me a drink with caffeine?
@@
CLUTRR
8
1,042
Q: [Carlos] is [Clarence]’s brother. [Carlos] and his sister, [Annie], went shopping. [Annie] asked her mom [Valerie] if she wanted anything, but [Valerie] said no. How is [Valerie] related to [Clarence]?
@@ E.2 Comparing Consistency Merics
Codex
.175†
.151
.159†
GPT-3.5-turbo
.205
.221†
.207
GPT-4
.116
.119†
.114
LLaMA-7B
.241†
.232
.235†
LLaMA-13B
.222†
.204
.211†
LLaMA-70B
.182†
.154
.165†
Mistral-7B
.205†
.183
.191†
Mistral-7B-instruct
.220†
.216
.215
Table 5: Overall Brier Score ($\downarrow$) of three consistency metrics averaged across all datasets and prompting strategies. † indicates that the current metric is significantly worse ($p<0.05$) than the best-performing metric (in bold).
@@ E.3 Calibration Results on All Datasets
ASDiv
.099
.090
.095
.205
.190
.150
.159
.120
GSM8K
.189
.158
.177
.252
.377
.262
.248
.188
Multi
.103
.085
.089
.187
.162
.135
.106
.117
SVAMP
.126
.103
.114
.193
.173
.139
.145
.148
Sport
.071
.068
.062
.346
.075
.067
.103
.200
Date
.174
.159
.162
.210
.251
.219
.191
.197
StrategyQA
.353
.213
.256
.305
.316
.257
.271
.220
CLUTRR
.288
.369
.327
.296
.536
.506
.330
.232
SayCan
.175
.115
.152
.243
.159
.145
.135
.190
average
.175
.151
.159
.249
.249
.209
.188
.179
@@
ASDiv
.194
.224
.194
.223
.213
n/a
n/a
n/a
GSM8K
.184
.196
.183
.260
.338
n/a
n/a
n/a
MultiArith
.044
.041
.039
.101
.065
n/a
n/a
n/a
SVAMP
.108
.115
.108
.164
.155
n/a
n/a
n/a
Sport
.089
.101
.095
.326
.151
n/a
n/a
n/a
Date
.266
.292
.280
.316
.348
n/a
n/a
n/a
StrategyQA
.393
.411
.376
.329
.342
n/a
n/a
n/a
CLUTRR
.429
.482
.465
.450
.509
n/a
n/a
n/a
SayCan
.137
.126 .126
.267
.341
n/a
n/a
n/a
average
.205
.221
.207
.271
.273
n/a
n/a
n/a
@@
ASDiv
.090
.103
.090
.091
.095
n/a
n/a
n/a
GSM8K
.083
.099
.087
.132
.144
n/a
n/a
n/a
MultiArith
.013
.010
.011
.015
.013
n/a
n/a
n/a
SVAMP
.047
.050
.047
.058
.063
n/a
n/a
n/a
Sport
.033
.031 .031
.160
.100
n/a
n/a
n/a
Date
.063
.069
.061
.073
.076
n/a
n/a
n/a
StrategyQA
.230
.205
.207
.195
.220
n/a
n/a
n/a
CLUTRR
.392
.443
.416
.386
.435
n/a
n/a
n/a
SayCan
.092
.065
.079
.279
.481
n/a
n/a
n/a
average
.116
.119
.114
.154
.181
n/a
n/a
n/a
Table 6: Brier Score ($\downarrow$) for closed-source LMs on all datasets, averaged across five prompting strategies. The best scores are in bold.
@@
ASDiv
.164
.155
.158
.411
GSM8K
.137
.166
.148
.713
MultiArith
.232
.231
.241
.544
SVAMP
.211
.195
.211
.455
Sport
.260
.221
.232
.272
Date
.216
.267
.235
.526
StrategyQA
.390
.265
.301
.408
CLUTRR
.290
.370
.323
.633
SayCan
.267
.214
.269
.307
average
.241
.232
.235
.474
@@
ASDiv
.144
.135
.136
.334
GSM8K
.177
.179
.181
.591
MultiArith
.232
.198
.222
.395
SVAMP
.205
.180
.200
.365
Sport
.170
.153
.151
.194
Date
.181
.206
.185
.402
StrategyQA
.383
.241
.285
.371
CLUTRR
.298
.353
.320
.596
SayCan
.209
.190
.220
.250
average
.222
.204
.211
.389
@@
ASDiv
.107
.099
.101
.209
GSM8K
.201
.168
.187
.375
MultiArith
.134
.108
.113
.188
SVAMP
.145
.112
.129
.183
Sport
.053
.041
.044
.046
Date
.167
.166
.168
.262
StrategyQA
.338
.192
.239
.289
CLUTRR
.287
.347
.309
.534
SayCan
.206
.156
.191
.179
average
.182
.154
.165
.252
@@
ASDiv
.122 .112
.113
.269
GSM8K
.197
.188
.196
.491
MultiArith
.189
.160
.171
.320
SVAMP
.176
.141
.162
.250
Sport
.116
.085
.090
.109
Date
.178
.202
.187
.373
StrategyQA
.363
.239
.276
.343
CLUTRR
.283
.334
.307
.555
SayCan
.225
.186
.217
.207
average
.205
.183
.191
.324
@@
ASDiv
.130
.127
.124
.288
GSM8K
.193
.191
.194
.508
MultiArith
.191
.164
.180
.319
SVAMP
.166
.147
.158
.274
Sport
.207
.196
.194
.222
Date
.220
.244
.227
.497
StrategyQA
.334
.269
.284
.366
CLUTRR
.306
.397
.347
.648
SayCan
.228
.209
.229
.332
average
.220
.216
.215
.384
Table 7: Brier Score ($\downarrow$) for closed-source LMs on all datasets, averaged across five prompting strategies. The best scores are in bold.
@@
MWP
.132
.077
.104
.237
.225
.156
.142
.108
MHQA
.188
.090
.119
.272
.214
.152
n/a
.167
Plan.
.203
.101
.159
.322
.159
.106
.117
.248
Relation.
.228
.368
.294
.214
.536
.512
.313
.175
average
.169
.117
.136
.256
.249
.189
.166
.151
@@
MWP
.118
.121
.115
.208
.193
n/a
n/a
n/a
MHQA
.230
.246
.233
.321
.277
n/a
n/a
n/a
Plan.
.154
.119
.128
.351
.357
n/a
n/a
n/a
Relation.
.426
.505
.471
.449
.519
n/a
n/a
n/a
average
.193
.205
.195
.289
.275
n/a
n/a
n/a
@@
MWP
.056
.064
.055
.053
.079
n/a
n/a
n/a
MHQA
.104
.089
.090
.139
.126
n/a
n/a
n/a
Plan.
.109
.061
.084
.351
.484
n/a
n/a
n/a
Relation.
.387
.454
.415
.381
.435
n/a
n/a
n/a
average
.114
.115
.110
.151
.179
n/a
n/a
n/a
@@
MWP
.138
.130
.141
.548
MHQA
.237
.189
.197
.400
Plan.
.256
.164
.259
.302
Relation.
.214
.359
.267
.641
average
.192
.179
.187
.482
@@
MWP
.159
.117
.151
.428
MHQA
.207
.142
.149
.319
Plan.
.205
.147
.217
.244
Relation.
.239
.334
.268
.602
average
.189
.153
.170
.391
@@
MWP
.153
.083
.120
.233
MHQA
.174
.082
.117
.193
Plan.
.229
.156
.207
.172
Relation.
.223
.333
.257
.539
average
.176
.118
.144
.247
@@
MWP
.165
.111
.138
.331
MHQA
.184
.120
.135
.271
Plan.
.260
.194
.231
.198
Relation.
.207
.295
.245
.561
average
.186
.144
.159
.322
@@
MWP
.154
.108
.127
.349
MHQA
.205
.191
.186
.355
Plan.
.203
.155
.212
.328
Relation.
.238
.402
.310
.659
average
.186
.174
.177
.383
Table 8: ECE score ($\downarrow$) for all LMs on four domains – Math Word Problems (MWP), Multi-hop QA (MHQA), Planning (Plan.), and Relational Inference (Relation.) – averaged across five prompting strategies. The best score is in bold.
Figure 9: ECE score ($\downarrow$) for each prompting strategy, averaged across all datasets and consistency metrics.

### v2

@@ Abstract
Accurately gauging the confidence level of Large Language Models’ (LLMs) predictions is pivotal for their reliable application. However, LLMs are often uncalibrated inherently and elude conventional calibration techniques due to their proprietary nature and massive scale. In this work, we explore the potential of deriving confidence from the distribution of multiple randomly sampled model generations, via three measures of consistency. We perform an extensive evaluation across various open and closed-source models on nine reasoning datasets. Results show that consistency-based calibration methods outperform existing post-hoc approaches. Meanwhile, we find that factors such as intermediate explanations, model scaling, and larger sample sizes enhance calibration, while instruction-tuning makes calibration more difficult. Moreover, confidence scores obtained from consistency have the potential to enhance model performance. Finally, we offer practical guidance on choosing suitable consistency metrics for calibration, tailored to the characteristics of various LMs. * Equal contribution; Qing Lyu did her work while interning at Allen Institute for AI. [^1] Code:https://github.com/veronica320/Calibrating-LLMs-with-Consistency
@@
Qing Lyu   Kumar Shridhar   Chaitanya Malaviya   Li Zhang   Yanai Elazar Niket Tandon   Marianna Apidianaki   Mrinmaya Sachan   Chris Callison-Burch University of Pennsylvania   ETH Zurich University of Washington   Allen Institute for AI lyuqing@sas.upenn.edushkumar@ethz.ch
@@ 3 Method
[…]
@@ 3.1 Calibration with Consistency
[…]
@@ 4 Experimental Setup
[…]
@@
[…]
@@
Codex .175 .151 .159 .249 .249 .209 .188 .179
GPT-3.5-turbo .205 .221† .207 .271 .273 n/a n/a n/a
GPT-4 .116 .119† .114 .154 .181 n/a n/a n/a
Table 1: Consistency metrics result in better Brier Scores than baselines ($\downarrow\downarrow$) for closed-source models. Scores are averaged across four domains and five prompting strategies. The best scores are in bold and the second-best scores are underlined. † indicates that the consistency metric performs statistically significantly better than the best-performing baseline ($p<0.05$).
@@
LLaMA-7B .241† .232 .235 .474
LLaMA-13B .222† .204 .211 .389
LLaMA-70B .182† .154 .165 .252
Mistral-7B .205† .183 .191 .324
Mistral-7B-instruct .220† .216 .215 .384
Table 2: Consistency metrics result in better Brier Scores ($\downarrow\downarrow$) than the logit baseline for open-source models.
@@
We consider the following LLMs: LLaMA (7B/13B/70B), Mistral (7B/7B-instruct), Codex, GPT-3.5-turbo, and GPT-4.555See checkpoint names and computational resources in Appendix B.
@@ 5.1 Comparing Calibration Methods
Figure 3: Brier Scores ($\downarrow\downarrow$) are improved with explanation-based prompting strategies, with Chain of Thought (CoT) and Faithful CoT (FCoT) performing the best. Scores here are averaged across all datasets and consistency metrics.
@@
Figure 4: The Brier Score ($\downarrow\downarrow$) tends to improve as the model size increases for the 3 studied calibration metrics across most of the prompting techniques we consider.
@@ 6.2 How Does Instruction-Tuning Affect Calibration?
Figure 5: Surprisingly, the non-instruction-tuned model (Mistral-7B) has better Brier Scores ($\downarrow\downarrow$) compared to an instruction-tuned model (Mistral-7B-instruct) across nearly all of our prompting strategies and tasks.
Figure 6: Brier Scores ($\downarrow\downarrow$) improve as we increase the number of samples for 3 of the 4 datasets. Results are obtained with GPT-3.5-turbo and CoT prompting.
Figure 7: Left: Consistency-based calibration outperforms the verbalized baselines in their ability to discriminate the correctness of predictions measured by Macro-F1 ($\uparrow\uparrow$). Right: Consistency-based calibration leads to a larger improvement in answer accuracy ($\uparrow\uparrow$) after correcting the top-k% most uncertain predictions with oracle answers. Scores in these figures are averaged across all prompting strategies.
@@ 6.3 How does the Number of Generated Outputs impact Calibration?
We analyze the usefulness of consistency-based calibration by generating different numbers of output samples and calculating different consistency metrics over them. Figure 6 demonstrates that generating more samples can lead to better calibration scores, as indicated by a downward trend in the Brier Score. We observe the improvement in the Brier Score as a function of the number of samples and the decision of the appropriate number of samples can be made based on the available computational budget and the desired calibration properties. Brier Scores usually saturate after 152015-20 samples, with a sharp drop at the beginning. For budget constraints, 353-5 samples is a good choice.
@@ 7 Case Study: Does Calibration Help Improve Model Performance?
The second experiment assesses the impact of a calibration method on answer accuracy, assuming an oracle subsequent self-correction mechanism. Given a model’s predictions $\hat{Y}$ on a dataset $X$, we identify the top-k% most uncertain predictions, $\hat{Y_{-}}$, which are those with the lowest confidence scores according to the calibration method, as incorrect. This fixed $k$ is chosen to be the true error rate of all model predictions, i.e., $k=1-\text{acc}(\hat{Y},X)$. Finally, we correct $\hat{Y_{-}}$ with the ground-truth answers and evaluate the resulting accuracy. As shown in Figure 7 (right), post-correction accuracy exceeds original accuracy to the greatest extent when applying consistency-based calibration.
@@ Appendix D Dataset Details
Math Word Problems GSM8K 8 1,319 Q: Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May?
@@
SVAMP 8 1,000 Q: Each pack of dvds costs 76 dollars. If there is a discount of 25 dollars on each pack. How much do you have to pay to buy each pack?
@@
MultiArith 8 600 Q: For Halloween Debby and her sister combined the candy they received. Debby had 32 pieces of candy while her sister had 42. If they ate 35 pieces the first night, how many pieces do they have left?
@@
ASDiv 8 2,096 Q: Seven red apples and two green apples are in the basket. How many apples are in the basket?
@@
StrategyQA 6 2,290 Q: Did Aristotle use a laptop?
@@
10 359 Q: Yesterday was April 30, 2021. What is the date tomorrow in MM/DD/YYYY?
@@
10 977 Q: Is the following sentence plausible: “Lebron James hit the turnaround jumper”?
@@
Planning SayCan 7 103 Q: Could you get me a drink with caffeine?
@@
CLUTRR 8 1,042 Q: [Carlos] is [Clarence]’s brother. [Carlos] and his sister, [Annie], went shopping. [Annie] asked her mom [Valerie] if she wanted anything, but [Valerie] said no. How is [Valerie] related to [Clarence]?
@@ E.2 Comparing Consistency Merics
Codex .175† .151 .159†
GPT-3.5-turbo .205 .221† .207
GPT-4 .116 .119† .114
LLaMA-7B .241† .232 .235†
LLaMA-13B .222† .204 .211†
LLaMA-70B .182† .154 .165†
Mistral-7B .205† .183 .191†
Mistral-7B-instruct .220† .216 .215
Table 5: Overall Brier Score ($\downarrow\downarrow$) of three consistency metrics averaged across all datasets and prompting strategies. † indicates that the current metric is significantly worse ($p<0.05$) than the best-performing metric (in bold).
@@ E.3 Calibration Results on All Datasets
ASDiv .099 .090 .095 .205 .190 .150 .159 .120
GSM8K .189 .158 .177 .252 .377 .262 .248 .188
Multi .103 .085 .089 .187 .162 .135 .106 .117
SVAMP .126 .103 .114 .193 .173 .139 .145 .148
Sport .071 .068 .062 .346 .075 .067 .103 .200
Date .174 .159 .162 .210 .251 .219 .191 .197
StrategyQA .353 .213 .256 .305 .316 .257 .271 .220
CLUTRR .288 .369 .327 .296 .536 .506 .330 .232
SayCan .175 .115 .152 .243 .159 .145 .135 .190
average .175 .151 .159 .249 .249 .209 .188 .179
@@
ASDiv .194 .224 .194 .223 .213 n/a n/a n/a
GSM8K .184 .196 .183 .260 .338 n/a n/a n/a
MultiArith .044 .041 .039 .101 .065 n/a n/a n/a
SVAMP .108 .115 .108 .164 .155 n/a n/a n/a
Sport .089 .101 .095 .326 .151 n/a n/a n/a
Date .266 .292 .280 .316 .348 n/a n/a n/a
StrategyQA .393 .411 .376 .329 .342 n/a n/a n/a
CLUTRR .429 .482 .465 .450 .509 n/a n/a n/a
SayCan .137 .126 .126 .267 .341 n/a n/a n/a
average .205 .221 .207 .271 .273 n/a n/a n/a
@@
ASDiv .090 .103 .090 .091 .095 n/a n/a n/a
GSM8K .083 .099 .087 .132 .144 n/a n/a n/a
MultiArith .013 .010 .011 .015 .013 n/a n/a n/a
SVAMP .047 .050 .047 .058 .063 n/a n/a n/a
Sport .033 .031 .031 .160 .100 n/a n/a n/a
Date .063 .069 .061 .073 .076 n/a n/a n/a
StrategyQA .230 .205 .207 .195 .220 n/a n/a n/a
CLUTRR .392 .443 .416 .386 .435 n/a n/a n/a
SayCan .092 .065 .079 .279 .481 n/a n/a n/a
average .116 .119 .114 .154 .181 n/a n/a n/a
Table 6: Brier Score ($\downarrow\downarrow$) for closed-source LMs on all datasets, averaged across five prompting strategies. The best scores are in bold.
@@
ASDiv .164 .155 .158 .411
GSM8K .137 .166 .148 .713
MultiArith .232 .231 .241 .544
SVAMP .211 .195 .211 .455
Sport .260 .221 .232 .272
Date .216 .267 .235 .526
StrategyQA .390 .265 .301 .408
CLUTRR .290 .370 .323 .633
SayCan .267 .214 .269 .307
average .241 .232 .235 .474
@@
ASDiv .144 .135 .136 .334
GSM8K .177 .179 .181 .591
MultiArith .232 .198 .222 .395
SVAMP .205 .180 .200 .365
Sport .170 .153 .151 .194
Date .181 .206 .185 .402
StrategyQA .383 .241 .285 .371
CLUTRR .298 .353 .320 .596
SayCan .209 .190 .220 .250
average .222 .204 .211 .389
@@
ASDiv .107 .099 .101 .209
GSM8K .201 .168 .187 .375
MultiArith .134 .108 .113 .188
SVAMP .145 .112 .129 .183
Sport .053 .041 .044 .046
Date .167 .166 .168 .262
StrategyQA .338 .192 .239 .289
CLUTRR .287 .347 .309 .534
SayCan .206 .156 .191 .179
average .182 .154 .165 .252
@@
ASDiv .122 .112 .113 .269
GSM8K .197 .188 .196 .491
MultiArith .189 .160 .171 .320
SVAMP .176 .141 .162 .250
Sport .116 .085 .090 .109
Date .178 .202 .187 .373
StrategyQA .363 .239 .276 .343
CLUTRR .283 .334 .307 .555
SayCan .225 .186 .217 .207
average .205 .183 .191 .324
@@
ASDiv .130 .127 .124 .288
GSM8K .193 .191 .194 .508
MultiArith .191 .164 .180 .319
SVAMP .166 .147 .158 .274
Sport .207 .196 .194 .222
Date .220 .244 .227 .497
StrategyQA .334 .269 .284 .366
CLUTRR .306 .397 .347 .648
SayCan .228 .209 .229 .332
average .220 .216 .215 .384
Table 7: Brier Score ($\downarrow\downarrow$) for closed-source LMs on all datasets, averaged across five prompting strategies. The best scores are in bold.
@@
MWP .132 .077 .104 .237 .225 .156 .142 .108
MHQA .188 .090 .119 .272 .214 .152 n/a .167
Plan. .203 .101 .159 .322 .159 .106 .117 .248
Relation. .228 .368 .294 .214 .536 .512 .313 .175
average .169 .117 .136 .256 .249 .189 .166 .151
@@
MWP .118 .121 .115 .208 .193 n/a n/a n/a
MHQA .230 .246 .233 .321 .277 n/a n/a n/a
Plan. .154 .119 .128 .351 .357 n/a n/a n/a
Relation. .426 .505 .471 .449 .519 n/a n/a n/a
average .193 .205 .195 .289 .275 n/a n/a n/a
@@
MWP .056 .064 .055 .053 .079 n/a n/a n/a
MHQA .104 .089 .090 .139 .126 n/a n/a n/a
Plan. .109 .061 .084 .351 .484 n/a n/a n/a
Relation. .387 .454 .415 .381 .435 n/a n/a n/a
average .114 .115 .110 .151 .179 n/a n/a n/a
@@
MWP .138 .130 .141 - - .548 - -
MHQA .237 .189 .197 - - .400 - -
Plan. .256 .164 .259 - - .302 - -
Relation. .214 .359 .267 - - .641 - -
average .192 .179 .187 - - .482 - -
@@
MWP .159 .117 .151 - - .428 - -
MHQA .207 .142 .149 - - .319 - -
Plan. .205 .147 .217 - - .244 - -
Relation. .239 .334 .268 - - .602 - -
average .189 .153 .170 - - .391 - -
@@
MWP .153 .083 .120 - - .233 - -
MHQA .174 .082 .117 - - .193 - -
Plan. .229 .156 .207 - - .172 - -
Relation. .223 .333 .257 - - .539 - -
average .176 .118 .144 - - .247 - -
@@
MWP .165 .111 .138 - - .331 - -
MHQA .184 .120 .135 - - .271 - -
Plan. .260 .194 .231 - - .198 - -
Relation. .207 .295 .245 - - .561 - -
average .186 .144 .159 - - .322 - -
@@
MWP .154 .108 .127 - - .349 - -
MHQA .205 .191 .186 - - .355 - -
Plan. .203 .155 .212 - - .328 - -
Relation. .238 .402 .310 - - .659 - -
average .186 .174 .177 - - .383 - -
Table 8: ECE score ($\downarrow\downarrow$) for all LMs on four domains – Math Word Problems (MWP), Multi-hop QA (MHQA), Planning (Plan.), and Relational Inference (Relation.) – averaged across five prompting strategies. The best score is in bold.
Figure 9: ECE score ($\downarrow\downarrow$) for each prompting strategy, averaged across all datasets and consistency metrics.